OpenStudy (anonymous):

Find all possible roots of the complex number [125/2(1-sq(3j)]^1/3

OpenStudy (anonymous):

no I will try and re-write the problem

zepdrix (zepdrix):

\[\Large\bf\sf \left[\frac{125}{2}(1-\sqrt3 \mathcal i\right]^{1/3}\]

zepdrix (zepdrix):

Like that maybe? :o

zepdrix (zepdrix):

Darn I missed a bracket :3

OpenStudy (anonymous):

\[[\frac{ 125 }{ 2 }(1-\sqrt{3}j)]^{\frac{ 1 }{ 3 }}\]

zepdrix (zepdrix):

\[\Large\bf\sf =\quad \left(\frac{125}{2}\right)^{1/3}\left(\color{royalblue}{1-\sqrt3 \mathcal j}\right)^{1/3}\]Focus on the blue part for a moment. We can rewrite these as trig functions of a specific angle.

zepdrix (zepdrix):

Can you think of what angle that would be?

zepdrix (zepdrix):

\[\Large\bf\sf \tan \theta\quad=\quad \frac{-\sqrt3}{1}\]

zepdrix (zepdrix):

The real component is `positive` the imaginary component is `negative` so we're in the ... 4th quadrant i guess.

OpenStudy (anonymous):

Yes Im following that ok

OpenStudy (anonymous):

where did the \[\frac{ 5\pi }{ 3 }\] come from?

zepdrix (zepdrix):

Oh crap I have the wrong angle... let's back up! :( I'll try to explain that better.

zepdrix (zepdrix):

Lemme erase some junk a minute.

zepdrix (zepdrix):

\[\Large\bf\sf =\quad \left(\frac{125}{2}\right)^{1/3}\left(\color{royalblue}{1-\sqrt3 \mathcal j}\right)^{1/3}\]We would like to turn the blue stuff into the special values that we know from the unit circle. If we can do that, we can convert these to sine and cosine. I don't like the thing I did with tangent, I think that was a little confusing. Let's try this instead. If we factor a 2 out of each term,\[\Large\bf\sf =\quad \left(\frac{125}{2}\right)^{1/3}\;2^{1/3}\left(\color{royalblue}{\frac{1}{2}-\frac{\sqrt3}{2} \mathcal j}\right)^{1/3}\]

zepdrix (zepdrix):

Ok maybe I did have the correct angle before lol. But uhh maybe this will make more sense... I hope at least -_-

zepdrix (zepdrix):

Do you understand what I did with the 2?

OpenStudy (anonymous):

Yes got that

zepdrix (zepdrix):

So then we just need to recall at what angle: cosine produces 1/2 sine produces -sqrt3/2

zepdrix (zepdrix):